# Rethinking Robustness in Machine Learning: A Posterior Agreement Approach

### v1

@@ 3 Methodology
In this section, we devise a tractable version of the posterior agreement for the validation of classification models in covariate shift settings. Similar to Buhmann et al. (2018), we start from the work of Rose (1998) for deterministic annealing, and adapt it in the context of PA to a classification task.
@@ 3.1 Setting
Let $D_{X}=(x_{i})_{i=1}^{N}$ be a dataset [^2] Note that, we do not need to specify the targets of $D_{X}$ for assessing the robustness of a model. of i.i.d. measurements (observations) drawn from a random variable $X$ with support $\mathcal{X}$. A $K$-class classifier can be defined as the composition of two functions:
  • a function $F:\mathcal{X}\rightarrow\mathbb{R}^{K}$, mapping observations to a vector of discriminant functions, i.e., $x\mapsto F(x)=(F_{1}(x),\dots,F_{K}(x))$. The discriminant functions are parametrized by a set of parameters $\theta$, which is fixed, as we are conducting an evaluation.
  • a decision rule $f:\mathbb{R}^{K}\rightarrow\{1,\dots,K\}$, applied to choose the class for each observation, based on the discriminant functions. Usually, $f$ is the Maximum A Posteriori (MAP) rule:
@@
Note that, since $F$ is fixed, the hypothesis set is spanned by all possible decision rules over $\mathcal{X}$. Restricting the process to $D_{X}$, gives rise to a finite hypothesis set [^3] The finiteness of $\mathcal{C}$ is a mathematical expedient to efficiently deal with the subsequent derivations. (cf. the discussion section). $\mathcal{C}$, containing all possible mappings from $D_{X}$ to $\{1,\dots,K\}$. Therefore, $|\mathcal{C}|=K^{N}$.
@@
$R(c,X;\theta)=\sum_{i=1}^{N}F_{c(x_{i})}(x_{i};\theta).$ (2)
Such score can be used to define a posterior $p(c\mid X)$ over the hypothesis set $\mathcal{C}$. In particular, we aim to find a probability distribution $p$ such that $R(c_{1},X;\theta)>R(c_{2},X;\theta)\iff p(c_{1}\mid X)>p(c_{2}\mid X)$, $\forall c_{1},c_{2}\in\mathcal{C}$, a property that we will rewrite as $\phi_{\mathcal{C}}(p)$. As many distributions fulfill this requirement, we narrow the possibilities by applying the Maximum Entropy Principle (MEP) (Jaynes, 1957), which states to search for the maximally uninformative (i.e., entropy-maximizing) distribution which best encodes the observed data.
@@
$\frac{\partial\mathcal{L}}{\partial p(c\mid X)}=-1-\log p(c\mid X)-\alpha+\beta R(c,X).$ (6)
Equating it to 00 and solving for $p(c)$ gives
$p(c\mid X)=\frac{\exp(\beta R(c,X))}{\exp(1+\alpha)}.$ (7)
Setting $\exp(1+\alpha)=\sum_{c\in\mathcal{C}}\exp(\beta R(c,X))$ and $\beta\geq 0$ ensures that $p(c\mid X)$ is a Gibbs distribution that satisfies the constraints 3.1 and 3.1 with the inverse temperature parameter $\beta$. The exact value of $\beta$ depends on $\mu$, and can be found by enforcing the posterior agreement principle, discussed in the next section.
@@
The energy term $F_{k}(x_{i})\in\mathbb{R}$ quantifies the degree of membership of observation $i$ to a class $k$ (the higher, the more probable). $F_{k}(x_{i})$ can be, for example, the logit output of a multi-class prediction model (e.g., a neural network).
@@ 3.2 Posterior Agreement
We are given two i.i.d. datasets $X^{\prime}$, $X^{\prime\!\>\!\prime}$ with $|X^{\prime}|=|X^{\prime\!\>\!\prime}|=N$ and for $i=\{1,\dots,N\}$ $x_{i}^{\prime}$ and $x_{i}^{\prime\!\>\!\prime}$ are two realizations sampled from the same ideal (i.e., noiseless) process $x_{i}^{(0)}$. Given a classification model, we can estimate its agreement in the prediction results through the expected posterior agreement kernel
@@
In the following, we propose and discuss an operative formula to compute the empirical posterior agreement kernel.
@@
By optimizing Program 3.2 with this kernel, we can see the effect of $\beta$ in determining the optimal overlap between posteriors. An increment in $\beta$ corresponds in peaking the distributions $p_{i}(\cdot\mid X)$, $X=X^{\prime},X^{\prime\!\>\!\prime}$ toward their MAP, awarding matching distributions that share the same MAP. On the contrary, decreasing $\beta$ flattens the distributions, and mitigates penalizations due to mismatching distributions with different MAPs. Therefore, the optimization operates a tradeoff between these opposite sub-objectives, searching for the best solution that explains the agreement between the two data sources.
We conclude by characterizing the posterior agreement metric as computed with the kernel of Theorem 2.
@@
Under no prior information available, PA has the following properties.
@@ 4 Experimental Results
In this section, we present a comprehensive analysis of the empirical results. In particular, we study two covariate shift scenarios, adversarial learning and domain generalization, and set up a comparison of the scores obtained by different models, using PA and accuracy-based metrics. Our main purpose is to highlight the differences between performance- and robustness-based evaluation criteria. Therefore, we compare several defences and learning techniques with different capabilities, to encompass, as much as possible, the variety of cases that can arise during an evaluation process. For further technical information, the reader is referred to our code implementation [^4] Metric: https://github.com/viictorjimenezzz/pa-metric
Experiments: https://github.com/viictorjimenezzz/pa-covariate-shift
For measuring model robustness, we use the PA metric with no prior information. In particular, we use the original test set as $X^{\prime}$ and its perturbed version as $X^{\prime\!\>\!\prime}$. For a better visualization of the results, we adopt a logarithmic scale, and we omit the $|\mathcal{C}|$ and the $1/N$ constants in PA to increase its dynamic range, therefore $\text{PA}\in[-N\log(K),0]$.
@@ 4.1 Adversarial Robustness
For the adversarial robustness scenarios, we carry out our experiments with the CIFAR-$10$ (Krizhevsky et al., 2009) dataset, widely adopted in the machine learning security literature as a benchmark for robustness evaluation. The dataset contains 60 0006000060\,00060 000 colour images of $32\times 32$ pixels equally distributed in $10$ classes. The analyzed models are trained on the training set (50 0005000050\,00050 000 images), and the PA evaluation is performed on the test set (10 0001000010\,00010 000 images). In particular, we perturb the test set with two different evasion attack methods: Projected Gradient Descent (PGD) (Madry et al., 2018) and Fast Minimum Norm (FMN) (Pintor et al., 2021). The max attack power is specified in terms of the $\ell_{\infty}$ norm and is set in advance ($\ell_{\infty}=[8/255,16/255,32/255]$) for the PGD attack, while it is automatically searched for each observation with the FMN attack. Both attacks are run for $1000$ steps on the test set to generate its perturbed version for subsequent PA computation.
We consider several models with different robustness properties. In particular, we score an undefended WideResNet-$28$-$10$ (Zagoruyko & Komodakis, 2016) and five defended models: a ResNet18181818 (He et al., 2016a) defended by Addepalli et al. (2022), a ResNet50505050 (He et al., 2016a) defended by Engstrom et al. (2019), a WideResNet-$28$-$10$ defended by Wang et al. (2023c), a PreActResNet18181818 (He et al., 2016b) defended by Wong et al. (2020) and a $3$-layer CNN (LeCun et al., 2015) defended by a JPEG pre-processing Das et al. (2017) and a Reverse Sigmoid post-processing Lee et al. (2018). All models except for the latter are implemented with the RobustBench library (Croce et al., 2021).
We perform a comparison of PA with an accuracy-based metric, the Attack Failure Rate ($\text{AFR}_{T}=1-\text{ASR}_{T}$) and evaluate the performance of the models under investigation to verify the effectiveness of the PGD and FMN attacks. The evaluation is performed over ten different adversarial ratios (ARs) of attacked examples in the test set, ordered by increasing attack power. In particular, a dataset with $AR=p$ contains the attacked observations under the $10p$-decile, ordered by $\ell_{\infty}$, and the remaining unattacked ones. $AR=1$ corresponds to an entirely attacked dataset. The $\beta$ parameter is searched with an Adam (Kingma & Ba, 2015) optimization procedure, run for $500$ epochs.
@@
In conclusion, while both metrics report similar trends and rankings, AFR tends to overestimate the robustness of the undefended model, especially when a few number of samples have been attacked, while PA provides a more consistent assessment, unbiased by the nature of the experiment (e.g., $\ell_{\infty}$ and $AR$).
Additionally, note that for $AR=0$, PA scores do not converge to the theoretical result of 00. In the absence of shift, the optimal $\beta$ tends toward infinity to peak all coinciding distributions over their MAPs (cf. proof of Theorem 3, Property $2$). However, the Adam optimizer progressively scales down the learning rate, dampening, in turn, the convergence rate. The obtained PA scores then reflect the “peakedness” of the initial model predictions, that is, they provide information about the predictive confidence of each model before the attack, with the most confident models scoring better. The ranking of PA coincides with that of $\text{AFR}_{T}$, hinting that in the absence of shift, the most confident models are also the more accurate [^5] However, this is not always necessarily true. We want to remark that task performance and predictive confidence are two distinct aspects (cf. Appendix C)..
@@ 4.2 Domain Generalization
Following the analysis on adversarial robustness, we perform a series of experiments on several domain generalization settings to assess both model discriminability and model selection when comparing PA to accuracy-based metrics. Again, $\text{AFR}_{T}$ is taken as the comparison reference. In addition, Adam is run for $1000$ epochs to search for the $\beta$ parameter.
In this scenario, we conduct our experiments through a modified version of the DiagViB-$6$ dataset (Eulig et al., 2021) that comprises distorted and upsampled coloured images of size $128\times 128$ from the MNIST dataset (LeCun, 1998). Each specific image factor is controllable, leading to changes in texture, hue, lightness, position, and scale, and observations paired to assemble $X^{\prime}$ and $X^{\prime\!\>\!\prime}$. Furthermore, we control the specific selected samples, differently from the common setting. In particular, we establish two unique domains for the training data $\{e_{0}^{1},e_{0}^{2}\}$ under a pair of original instantiations of each image factor, where all factors from $e_{0}^{1}$ are different from $e_{0}^{2}$. We induce a second pair of domains for the validation data, following precisely the factor configuration of the training domains. Finally, a set of six additional domains, $\{e_{0},e_{1},\dots,e_{5}\}$, is generated, with each $e_{i}$ corresponding to images where $i$ factors have been modified with respect to $e_{0}^{1}$. Our final dataset comprises two sets of 40 0004000040\,00040 000 images for training, two sets of 20 0002000020\,00020 000 images for validation, and six sets of 10 0001000010\,00010 000 images for testing. Note that the original MNIST images used for generating the datasets do not overlap. For more details on how the dataset was adjusted to our setting, refer to Appendix E.
For the experiment, we consider a ResNet50505050 model trained with three different algorithms: vanilla Empirical Risk Minimization (ERM), Invariant Risk Minimization (IRM) (Arjovsky et al., 2019) and Learning Invariant Predictors with Selective Augmentation (LISA) (Yao et al., 2022).
@@
Figure 8: Average predictive confidence for three models—(blue) Vanilla ERM, (orange) IRM, and (green) LISA, under a $SF=1$. Each panel shows results at a different number of shift factors (#$\#SF$), illustrating how the models’ predictive confidence varies as the distribution shift increases. The left column remains fixed, and the right column changes across #$\#SF=1,2,3,4,5$. The final bar plot on the right depicts the $\beta$ for $5$ shift factors. We see that higher model confidence leads to an adjusted higher $\beta$.
As depicted, and already well established in Koh et al. (2021), the overall trend of robustness starts with ERM as the weakest model, followed by IRM, and then by LISA as the most robust. However, PA is also able to identify more nuanced robustness behaviors. In particular, when $SF=1$, ERM achieves a higher PA score because it relies on domain-specific features already encountered during training, even when the number of mismatching predictions is higher. PA is the only metric that was able to discriminate ERM. Conversely, the domain invariance imposed by IRM slightly hinders PA under these mild shifts, even though its overall AFR remains comparable or better.
This subtle source-versus-target domain difference clearly illustrates why PA is more suitable than $\text{AFR}_{T}$ for evaluating robustness: the former captures how models respond to domain shifts that partially overlap with training distributions, whereas the latter might mask these fine-grained effects. Consequently, IRM’s robust design trades off some PA in low-shift regimes but does not sacrifice performance, distinguishing its behavior from both ERM and LISA in a way that PA can clearly reveal.
@@
We also attempted to better understand the impact of controlled domain shifts and their impact on robustness assessment through PA (cf. Appendix D for more details). A phenomenon observed in this work is the decrease in classification performance as the number of shift factors increases, while the PA metric remains comparable or slightly improves. Specifically, at $SF=5$, the feature space differs significantly (as shown in Figure 11) from those at $SF\leq 4$, leading to lower accuracy yet roughly similar (or slightly improved) robustness. This observation highlights that in more complex settings, where the effects of the covariate shift are directly controlled (as in the case of DiagViB-$6$), the robustness of the model does not stem solely from how out-of-distribution the samples are. Instead, as noted by Geirhos et al. (2020), it is influenced by compounded learning phenomena related to which features are captured during training. Consequently, future research should account for these interactions when designing evaluation protocols. For example, one possibility is to test the model on all possible combinations of shift factors, thereby more comprehensively revealing how different combinations and sequences of shifts affect robustness.

### v2

@@ 3 Methodology
Figure 2: The PA framework, applied in the estimation of a real-valued parameter (e.g., a distribution’s mean). Two posteriors $p(c\mid{X}^{{\prime}})$, $p(c\mid{X}^{\prime\!\>\!\prime})$, $c\in\mathbb{R}$ are fit to two different data realizations. The optimization over the inverse temperature parameter $\beta$ is required to make the two posteriors insensitive to sampling noise. When $\beta$ is too low (top left), the posteriors tend to be uniform and uninformative, i.e., they underfit. When $\beta$ is too high (top right), the posteriors are too peaked and therefore sensitive to noise perturbation, i.e., they overfit. An optimal $\beta$ (top center) maximizes the posterior agreement (bottom), ensuring informativeness and stability that can be then used for model selection or, in our case, for robustness assessment. Illustration based on Buhmann et al. (2018a).
In this section, we devise a tractable version of the posterior agreement for the validation of classification models in covariate shift settings. We first introduce the original PA framework and then adapt it in the context of a classification task. Similar to Buhmann (2010), we take our inspiration from simulated annealing and its deterministic variants (Rose et al., 1992; Buhmann & Kühnel, 1993; Rose, 1998). Our notation follows Buhmann et al. (2018a).
3.1 Background
In the original posterior agreement framework, we are given two datasets $^{{\prime}}, X′′𝒳$$^{\prime\!\>\!\prime}\in\mathcal{X}$, derived from the same experimental settings, thus, with identical signal, but different noise realizations. In classic optimization settings, both data instances can be used separately to train a model, i.e., to select a solution $c$ from a hypothesis set $\mathcal{C}$. Such solutions are, in general, distinct and often explain not only the signal but also part of the noise. To counter this overfitting, the PA approach (Figure 2) focuses on modelling two posteriors $p(c\mid{X}^{{\prime}})$, $p(c\mid{X}^{\prime\!\>\!\prime})$. In particular, PA mitigates the effect of the sampling noise by modulating the dispersion of the distributions. Large dispersion of a flat posterior can explain many different data samples and, therefore, they are less informative. On the other hand, small dispersion of a peaked posterior is highly informative but does not generalize well over sampling noise, and is, thus, less stable. PA aims at balancing these two aspects, informativeness and stability, to achieve an optimal agreement between the posteriors that explains the selected solutions. The obtained distributions can then be combined and used to sample a more robust solution that better generalizes over the two noisy data realizations.
Traditionally, posteriors have been designed by calculating or approximating the respective Gibbs distribution defined by the score function $\mathcal{S}(c,X)$ used in the optimization problem. In PA, both Gibbs distributions are parameterized by a shared inverse computational temperature $\beta$ controlling their dispersion and their agreement. An optimal control of $\beta$ renders the posteriors informative and insensitive to sampling noise perturbation (cf. Figure 2, top).
In our case, $^{{\prime}}, X′′{X}^{\prime\!\>\!\prime}$ represent a test set and its shifted instance, while the Gibbs posteriors stem from the model predictions of a trained model. The optimal agreement can then be quantified and used as a measure to validate the robustness against covariate shift. A robust model will generate similar posteriors, thus a high agreement. In a brittle model with lack of robustness, the posteriors will differ substantially.
@@ 3.2 Setting
Let $D_{X}=(x_{i})_{i=1}^{N}$ be a dataset [^2] Note that we do not need to specify the targets of $D_{X}$ for assessing the robustness of a model. of i.i.d. measurements (observations) drawn from a random variable $X$ with support $\mathcal{X}$. A $K$-class classifier can be defined as the composition of two functions:
  • a function $F:\mathcal{X}\rightarrow\mathbb{R}^{K}$, mapping observations to a tuple of class conditional score functions, i.e., $x\mapsto F(x)=(F_{1}(x),\dots,F_{K}(x))$. These functions are parametrized by a set of parameters $\theta$, which is fixed, as we are conducting an evaluation. The score term $F_{k}(x)\in\mathbb{R}$ quantifies the membership degree of observation $x$ to class $k$ (the higher, the more probable). $F_{k}(x)$ can be, for example, the logit of a multi-class prediction model (e.g., a neural network).
  • a decision rule $f:\mathbb{R}^{K}\rightarrow\{1,\dots,K\}$, applied to choose the class for each observation, based on the class conditional score functions. Usually, $f$ is the Maximum A Posteriori (MAP) rule:
@@
Note that, since $F$ is fixed, the hypothesis set is spanned by all possible decision rules over $\mathcal{X}$. Restricting the process to $D_{X}$, gives rise to a finite hypothesis set $\mathcal{C}$, containing all possible mappings from $c:D_{X}\rightarrow\{1,\dots K\}$, induced by a specific classifier $c\in\mathcal{C}$  [^3] The finiteness of $\mathcal{C}$ is a mathematical expedient to efficiently deal with the subsequent derivations. By restricting to $D_{X}$, $c$ represents an index mapping on a finite set, which is less discriminative than $c=f\circ F$ defined above. cf. also Section 5.. Therefore, $|\mathcal{C}|=K^{N}$.
@@
$\mathcal{S}(c,X;\theta)=\sum_{i=1}^{N}F_{c(x_{i})}(x_{i};\theta).$ (2)
Such a score can be used to define a posterior $p(c\mid X)$ over the hypothesis set $\mathcal{C}$. In particular, we aim to find a probability distribution $p$ such that $\mathcal{S}(c_{1},X;\theta)>\mathcal{S}(c_{2},X;\theta)\iff p(c_{1}\mid X)>p(c_{2}\mid X)$, $\forall c_{1},c_{2}\in\mathcal{C}$, a property that we will rewrite as $\phi_{\mathcal{C}}(p)$. As many distributions fulfill this requirement, we narrow the possibilities by applying the Maximum Entropy Principle (MEP) (Jaynes, 1957), which states to search for the maximally uninformative (i.e., entropy-maximizing) distribution which best encodes the observed data.
The following optimization objective
@@
[…]
$p(c\mid X)$ $\geq 0,$ (3c)
$\sum_{c\in\mathcal{C}}p(c\mid X)$ $=1,$ (3d)
$\mathbb{E}_{c\,\mid\,X}[\mathcal{S}(c,X)]$ $=\mu$ (3e)
gives the required solution. Here, $\mu$ is a hyperparameter ensuring that the expectation is finite. We set the Lagrangian without the inequality constraints
@@
$\frac{\partial\mathcal{L}}{\partial p(c\mid X)}=-1-\log p(c\mid X)-\alpha+\beta\mathcal{S}(c,X).$ (7)
Equating it to 0 and solving for $p(c)$ gives
$p(c\mid X)=\frac{\exp(\beta\mathcal{S}(c,X))}{\exp(1+\alpha)}.$ (8)
Setting $\exp(1+\alpha)=\sum_{c\in\mathcal{C}}\exp(\beta\mathcal{S}(c,X))$ and $\beta\geq 0$ ensures that $p(c\mid X)$ is a Gibbs distribution that satisfies the constraints 3b and 3c with the inverse temperature parameter $\beta$. The exact value of $\beta$ depends on $\mu$, and can be found by enforcing the posterior agreement principle, discussed in the next section.
@@ 3.3 Posterior Agreement
We are given two i.i.d. datasets $^{\prime}, X′′X^{\prime\!\>\!\prime}$ with $|X^{\prime}|=|X^{\prime\!\>\!\prime}|=N$ and for $i=\{1,\dots,N\}$ $x_{i}^{\prime}$ and $x_{i}^{\prime\!\>\!\prime}$ are two realizations sampled from the same ideal (i.e., noiseless) process $x_{i}^{(0)}$. Given a classification model, we can estimate its agreement in the prediction results through the expected posterior agreement kernel
@@
The computation depends on the inverse temperature hyperparameter $\beta\in\mathbb{R}_{\geq 0}$. Similar to the MEP, we search for the $\beta$ that maximizes the overlap between the posteriors. The Posterior Agreement (PA) measure is then computed as
@@
[…]
$1/N$ is a scaling correction factor since PA increases proportionally with the size of the dataset.
Below, we propose and discuss an operative formula to compute the empirical posterior agreement kernel.
@@
By optimizing Program 13 with this kernel, we can see the effect of $\beta$ in determining the optimal overlap between posteriors. An increment in $\beta$ corresponds in peaking the distributions $p_{i}(\cdot\mid X)$, $X=X^{\prime},X^{\prime\!\>\!\prime}$ toward their MAP, awarding matching distributions that share the same MAP. On the contrary, decreasing $\beta$ flattens the distributions, and mitigates penalizations due to mismatching distributions with different MAPs. Therefore, the optimization operates a tradeoff between these opposite sub-objectives, searching for the best solution that explains the agreement between the two data sources.
We conclude by characterizing the posterior agreement measure as computed with the kernel of Theorem 2.
@@
Under no prior information available, the PA measure has the following properties:
@@ 4 Experimental Results
In this section, we present a comprehensive analysis of the empirical results. In particular, we study two covariate shift scenarios, adversarial learning and domain generalization, and set up a comparison of the scores obtained by different models, using PA and accuracy-based measures. Our main purpose is to highlight the differences between performance- and robustness-based evaluation criteria. Therefore, we compare several defences and learning techniques with different capabilities, to encompass, as much as possible, the variety of cases that can arise during an evaluation process. For further technical information, the reader is referred to our code implementation [^4] PA: https://github.com/viictorjimenezzz/pa-metric
Experiments: https://github.com/viictorjimenezzz/pa-covariate-shift.
For measuring model robustness, we use the PA measure with no prior information. In particular, we use the original test set as $X^{\prime}$ and its perturbed version as $X^{\prime\!\>\!\prime}$. For a better visualization of the results, we adopt a logarithmic scale, and we omit the $|\mathcal{C}|$ and the $1/N$ constants in PA to increase its dynamic range, therefore $\text{PA}\in[-N\log(K),0]$.
@@ 4.1 Adversarial Robustness
For the adversarial robustness scenarios, we carry out our experiments with the CIFAR-$10$ (Krizhevsky et al., 2009) and the ImageNet (Deng et al., 2009) datasets, widely adopted in the machine learning security literature as benchmarks for robustness evaluation. The CIFAR-$10$ dataset contains 60 00060\,000 colour images of $32\times 32$ pixels equally distributed in $10$ classes. The analyzed models are trained on the training set (50 00050\,000 images), and the PA evaluation is performed on the test set (10 00010\,000 images). The ImageNet dataset contains circa 1.51.5M colour images processed at $256\times 256$ pixels, unequally distributed in $1000$ classes. In this case, the evaluation is conducted on the validation set, as is usually done in the literature. A random subset of 10 00010\,000 images is selected, so that finite-sample effects and PA values are comparable to those in CIFAR-$10$ experiments. In both cases, we perturb the test set using several evasion attack methods. We start from two simple attacks, Projected Gradient Descent (PGD, Madry et al., 2018) and Fast Minimum Norm (FMN, Pintor et al., 2021), and we proceed to more efficient benchmark attacks such as APGD and FAB from the AutoAttack library (Croce & Hein, 2020b). The attack power is specified in terms of the $\ell_{\infty}$ norm and is set in advance ($\ell_{\infty}=\ell/225$, $\ell\in[2,4,8,16,32]$) for all attacks except for FMN, where the attack norm is automatically searched for each observation. All attacks are run on the evaluation set to produce perturbed versions for subsequent PA computation. PGD and FMN attacks are run for $1000$ steps, while the number of steps in AutoAttack is determined automatically. In the following, we discuss the results for PGD and FMN in CIFAR-$10$, while we include the remaining AutoAttack experiments in Appendix F.
We consider several models with different robustness properties. In particular, we score an undefended WideResNet-2828-$10$ (Zagoruyko & Komodakis, 2016) and five defended models: a ResNet1818 (He et al., 2016a) defended by Addepalli et al. (2022), a ResNet5050 (He et al., 2016a) defended by Engstrom et al. (2019), a WideResNet-2828-$10$ defended by Wang et al. (2023c), a PreActResNet1818 (He et al., 2016b) defended according to Wong et al. (2020) and a $3$-layer CNN (LeCun et al., 2015) defended by a JPEG pre-processing Das et al. (2017) and a Reverse Sigmoid post-processing Lee et al. (2018). All models except for the latter are implemented with the RobustBench library (Croce et al., 2021).
We perform a comparison of PA with an accuracy-based measure, the Attack Failure Rate ($\text{AFR}_{T}=1-\text{ASR}_{T}$) and evaluate the performance of the models under investigation to verify the effectiveness of the PGD and FMN attacks. The evaluation is performed over ten different adversarial ratios (ARs) of attacked examples in the test set, ordered by increasing attack power. In particular, a dataset with $AR=p$ contains the attacked observations under the $10p$-decile, ordered by $\ell_{\infty}$, and the remaining unattacked ones. $AR=1$ corresponds to an entirely attacked dataset. The $\beta$ parameter is searched with an Adam (Kingma & Ba, 2015) optimization procedure, run for $500$ epochs. We attested a relatively, fast optimization of the $\beta$ parameter, usually in the order of tens of minutes in single GPU even for large dataset (i.e., ImageNet).
@@
In conclusion, while both measures report similar trends and rankings, AFR tends to overestimate the robustness of the undefended model, especially when a few number of samples have been attacked, while PA provides a more consistent assessment, unbiased by the nature of the experiment (e.g., $\ell_{\infty}$ and $AR$).
Additionally, note that for $AR=0$, PA scores do not converge to the theoretical result of 0. In the absence of shift, the optimal $\beta$ tends toward infinity to peak all coinciding distributions over their MAPs (cf. proof of Theorem 3, Property $2$). However, the Adam optimizer progressively scales down the learning rate, dampening, in turn, the convergence rate. The obtained PA scores then reflect the “peakedness” of the initial model predictions, that is, they provide information about the predictive confidence of each model before the attack, with the most confident models scoring better. The ranking of PA coincides with that of $\text{AFR}_{T}$, hinting that in the absence of shift, the tested models yield higher confidence in positive predictions than in negative ones [^5] However, this is not always necessarily true. We want to remark that task performance and predictive confidence are two distinct aspects (cf. Appendix C).. This additional insight is obtained thanks to the use of an adaptive optimization algorithm such as Adam, advocating for its use in the optimization of $\beta$.
@@
Autoattack
In Appendix F, we present experiments conducted on CIFAR-$10$ and ImageNet with the Autoattack library (Croce & Hein, 2020b). Specifically, we report robustness scores for APGD-CE/DLR/DLR, and FAB attacks on CIFAR-$10$. The results exhibit trends consistent with those discussed above: PA provides a reliable assessment under increasing shift power, whereas $\text{AFR}_{T}$ yields inconsistent rankings and fails to discriminate effectively between models. Below, we shortly discuss the obtained results. The reader is referred to the appendix for a more detailed discussion.
In the CIFAR-$10$ case, Autoattack is much more effective than the vanilla PGD attack discussed above. Evaluating a wider range of epsilon values reveals the full spectrum of $\text{AFR}_{T}$ scores and confirms the superior discriminability of PA even when all defenses fail. For example, PA ranks Addepalli et al. (2022) and Wang et al. (2023c) as the most robust defenses against APGD attacks across all $\ell_{\infty}$ and AR values, whereas $\text{AFR}_{T}$ only reaches this conclusion in the high shift power regime.
In the ImageNet case, we observe that PA and $\text{AFR}_{T}$ yield the same defense ranking. This is due to the high success of the attack, which makes performance degradation drive the PA score. Still, PA demonstrates greater consistency, ranking models correctly when $AR>0$, whereas $\text{AFR}_{T}$ requires $AR\gtrsim 0.8$.
@@ 4.2 Domain Generalization
Following the analysis on adversarial robustness, we perform a series of experiments on several domain generalization settings to assess both model discriminability and model selection when comparing PA to accuracy-based measures. Again, $\text{AFR}_{T}$ is taken as the comparison reference. In addition, Adam is run for $1000$ epochs to search for the $\beta$ parameter.
In this scenario, we conduct our experiments through a modified version of the DiagViB-$6$ dataset (Eulig et al., 2021) that comprises distorted and upsampled coloured images of size $128\times 128$ from the MNIST dataset (LeCun, 1998). Each image is defined by five controllable factors that induce changes in texture, hue, lightness, position, and scale. A domain corresponds to a specific configuration of these factors.
For training, we define two unique original domains, $\{e_{0}^{1},e_{0}^{2}\}$, under a pair of original instantiations of each image factor, where all the factors of $e_{0}^{1}$ are different from $e_{0}^{2}$. We induce a second pair of domains for the validation data, following precisely the factor configuration of the training domains, but with disjoint set of samples. Finally, a set of six additional domains, $\{e_{0},e_{1},\dots,e_{5}\}$, is generated. The factors of domain $e_{0}$ are identical $e_{0}^{1}$, while each $e_{i}(i=1,\dotsc,5)$ is obtainined by modifying exacty $i$ of the five factors relative to $e_{0}^{1}$.
Our final dataset comprises two sets of 40 00040\,000 images for training, two sets of 20 00020\,000 images for validation, and six sets of 10 00010\,000 images for testing. Note that the original MNIST images used for generating the datasets do not overlap. For more details on how the dataset was adjusted to our setting, refer to Appendix E.
For the experiments, we consider a ResNet5050 model trained with three different algorithms: vanilla Empirical Risk Minimization (ERM, Vapnik, 1991), Invariant Risk Minimization (IRM, Arjovsky et al., 2019) and Learning Invariant Predictors with Selective Augmentation (LISA, Yao et al., 2022). These methods were chosen as representative examples of two distinct approaches to domain generalization: IRM operates by enforcing invariance at the level of latent representations, while LISA augments the data itself to promote invariance in the input space. By including both, we evaluate robustness across approaches that intervene at different stages of the learning process.
@@
Figure 8: Average predictive confidence for three models—(blue) Vanilla ERM, (orange) IRM, and (green) LISA, under a $SF=1$. Each panel shows results at a different number of shift factors (#SF\#SF), illustrating how the models’ predictive confidence varies as the distribution shift increases. The left column remains fixed, and the right column changes across #SF=1,2,3,4,5\#SF=1,2,3,4,5. The final bar plot on the right depicts the $\beta$ for 55 shift factors. We see that higher model confidence leads to an adjusted higher $\beta$.
As depicted, and already well established in Koh et al. (2021), the overall trend of robustness starts with ERM as the weakest model, followed by IRM, and then by LISA as the most robust. However, PA is also able to identify more nuanced robustness behaviors. In particular, when $SF=1$, ERM achieves a higher PA score because it relies on domain-specific features already encountered during training, even when the number of mismatching predictions is higher. Note that PA is the only measure that was able to discriminate ERM. Conversely, the domain invariance imposed by IRM slightly hinders PA under these mild shifts, even though its overall AFR remains comparable or better.
This illustrates that PA and $\text{AFR}_{T}$ provide complementary insights in domain generalization: AFR reflects performance differences across shifts, while PA emphasizes the stability of predictive distributions under factor perturbations. For example, PA stabilizes after certain shifts because the model has not learned to represent additional factors (e.g., after hue and brightness), whereas AFR continues to decrease as performance drops. At the same time, PA can capture fine-grained effects that $\text{AFR}_{T}$ may overlook. In particular, the source-versus-target domain difference highlights how PA reveals model behavior when domain shifts partially overlap with training distributions, while $\text{AFR}_{T}$ may mask these nuances. Consequently, IRM’s robust design trades off some PA in low-shift regimes but does not sacrifice performance, distinguishing its behavior from both ERM and LISA in a way that PA can clearly reveal. Taken together, both measures are valuable for a comprehensive evaluation of robustness.
@@
We also attempted to better understand the impact of controlled domain shifts and their impact on robustness assessment through PA (cf. Appendix D for more details). A phenomenon observed in this work is the decrease in classification performance as the number of shift factors increases, while the PA measure remains comparable or slightly improves. Specifically, at $SF=5$, the feature space differs significantly (as shown in Figure 11) from those at $SF\leq 4$, leading to lower accuracy yet roughly similar (or slightly improved) robustness. This observation highlights that in more complex settings, where the effects of the covariate shift are directly controlled (as in the case of DiagViB-$6$), the robustness of the model does not stem solely from how out-of-distribution the samples are. Instead, as noted by Geirhos et al. (2020), it is influenced by compounded learning phenomena related to which features are captured during training. Consequently, future research should account for these interactions when designing evaluation protocols. For example, one possibility is to test the model on all possible combinations of shift factors, thereby more comprehensively revealing how different combinations and sequences of shifts affect robustness.